# LLM Knows Geometry Better than Algebra: Numerical Understanding of LLM-Based Agents in A Trading Arena

### v1

@@ Acknowledgments
This work was supported in part by the National Natural Science Foundation of China under Grants 62301213 and 62271361, the Hubei Provincial Key Research and Development Program under Grant 2024BAB039, and the Open Project Funding of the Hubei Key Laboratory of Big Data Intelligent Analysis and Application, Hubei University under Grant 2024BDIAA01.
@@ Appendix A Agent Trading Arena
A.1 Agent Trading Arena Details
As illustrated in Figure 5, each agent’s workflow integrates LLMs for chat pool interactions, stock analysis, decision-making, and reflection. In the stock analysis and decision-making modules, all outputs are validated for consistency with both common sense and operational requirements before execution.
@@ A.1.1 Action Decision-Making
Action generation follows the LLM framework. The agent responsible for generating actions receives corresponding prompts via SQLite. Based on these prompts and specified output formats, the agent decides whether to buy, sell, or hold stocks. The action generation process is outlined in Equation 1, with input prompts shown in Figure 6 and the corresponding outputs displayed in the adjacent figure.
@@ A.1.2 Environmental Interaction
To isolate external influences, we created a virtual sandbox environment where each agent is assigned a unique ID, and their actions affect the environment. The function $\phi$ facilitates environmental interactions, as shown in Algorithm 1, where “OPS” retrieves agent actions, “date” refers to the trading date, and “Z” represents the memory used for interaction with the environment. Through $\phi$, each agent’s actions, such as buying or selling stocks, determine the current stock price and update the trading platform, including stock prices and available shares. Stock prices are independent of external factors and are influenced solely by the sandbox’s internal dynamics. The stock price is updated with each transaction according to the following formula:
@@
             if $<P.\text{Cash}$ Cash then
@@
[…]
@@ A.1.4 Reflection
Together, short-term and long-term memory provide essential context for the agents. Success in this environment depends on their ability to understand the game rules and develop strategies that outmaneuver competitors. Agents continuously refine their strategies based on incremental quantitative feedback, adjusting their actions to align with long-term objectives.
@@
Appendix B NASDAQ Stock
This study selects seven stocks from the NASDAQ exchange: AAPL, AMZA, GOOGL, MSFT, NFLX, NVDA, and TSLA. These stocks represent leading companies in the technology, energy, and automotive sectors, providing high market representativeness and significant trading volumes. The NASDAQ stock dataset from Yahoo Finance spans July 3, 2023, to October 29, 2024, excluding weekends and holidays. This dataset reflects current market trends and serves as a timely foundation for our research. It includes daily records of opening price, closing price, highest price, lowest price, and trading volume, as well as relevant technical indicators, offering a comprehensive view of market behavior.
The training and testing periods for MACD, StockFormer (Gao et al., 2023), TimesNet (Wu et al., 2023), and our system are shown in Table 7. StockFormer and TimesNet require longer training periods, with their training datasets spanning a broader time range compared to the other models. In contrast, our system does not require training and relies solely on historical stock data for making trading decisions.
Strategy Training Testing
Start End Start End
MACD 29/7/2024 30/8/2024 3/9/2024 29/10/2024
StockFormer 3/7/2023 30/8/2024 3/9/2024 29/10/2024
TimesNet 3/7/2023 30/8/2024 3/9/2024 29/10/2024
Ours 16/8/2024 30/8/2024 3/9/2024 29/10/2024
Table 7: Training and Testing Periods for Various Models on the NASDAQ Stock Dataset.
Appendix C Visualization Input
Figure 7 illustrates the system’s input prompts and corresponding outputs during the strategy update process. The input prompts consist of both textual and visual components, including daily K-line charts, transaction histories, and agent trading volumes, all of which inform the strategy update.
Appendix D Simulation Process
In the Agent Trading Arena, the simulation process unfolds as follows: First, rumors are generated in the chat pool based on the previous day’s stock market analysis. Next, historical stock data is analyzed, followed by decision-making and execution. Short-term memory is formed through interactions with the environment. Finally, the system evaluates this memory, updates the strategy, and consolidates it into long-term memory. This entire process is illustrated in Figure 8.
Appendix E Simulation Process
The experiments involved several LLMs, including LLaMa-3 (Dubey et al., 2024), GPT-4o (Hurst et al., 2024), DeepSeek (Liu et al., 2024), Qwen-2.5 (Yang et al., 2024b), and Gemini-1.5 (Reid et al., 2024), corresponding to the models Meta-LLaMa-3-70B-Instruct, gpt-4o-2024-08-06, DeepSeek-chat, Qwen2.5-72B-Instruct, and Gemini-1.5-pro.

### v2

@@ Acknowledgments
This work was supported in part by the National Natural Science Foundation of China under Grants 62301213 and 62271361, the Hubei Provincial Key Research and Development Program under Grant 2024BAB039, and the Open Project Funding of the Hubei Key Laboratory of Big Data Intelligent Analysis and Application, Hubei University, under Grant 2024BDIAA01.
@@ Appendix A Agent Trading Arena
A.1 Overview
As illustrated in Fig.˜7, each agent’s workflow integrates LLMs for chat pool interactions, stock analysis, decision-making, and reflection. In the stock analysis and decision-making modules, all outputs are validated for consistency with both common sense and operational requirements before execution.
@@ A.1.1 Action Decision-Making
Action generation follows the LLM framework. The agent responsible for generating actions receives corresponding prompts via SQLite. Based on these prompts and specified output formats, the agent decides whether to buy, sell, or hold stocks. The action generation process is outlined in Eq.˜1, with input prompts shown in Fig.˜8 and the corresponding outputs displayed in the adjacent figure.
@@ A.1.2 Environmental Interaction
To isolate external influences, we created a virtual sandbox environment where each agent is assigned a unique ID, and their actions affect the environment. The function $\phi$ facilitates environmental interactions, as shown in Algorithm˜1, where “OPS” retrieves agent actions, “date” refers to the trading date, and “Z” represents the memory used for interaction with the environment. Through $\phi$, each agent’s actions, such as buying or selling stocks, determine the current stock price and update the trading platform, including stock prices and available shares. Stock prices are independent of external factors and are influenced solely by the sandbox’s internal dynamics. The stock price is updated with each transaction according to the following formula:
@@
    if $\text{Cash}<P.\text{Cash}$ then
@@
$\text{Market}(\text{date},\text{Persons})$;
@@ A.1.4 Reflection
Together, short-term and long-term memory provides essential context for the agents. Success in this environment depends on their ability to understand the game rules and develop strategies that outmaneuver competitors. Agents continuously refine their strategies based on incremental quantitative feedback, adjusting their actions to align with long-term objectives.
Appendix B NASDAQ and CSI
This study selects seven stocks from the NASDAQ exchange: AAPL, AMZA, GOOGL, MSFT, NFLX, NVDA, and TSLA. These stocks represent leading companies in the technology, energy, and automotive sectors, providing high market representativeness and significant trading volumes. The NASDAQ dataset from Yahoo Finance spans July 3, 2023, to October 29, 2024, excluding weekends and holidays. This dataset reflects current market trends and serves as a timely foundation for our research. It includes daily records of opening price, closing price, highest price, lowest price, and trading volume, as well as relevant technical indicators, offering a comprehensive view of market behavior.
This study also selects seven representative stocks from the Chinese stock market (CSI), covering various sectors with industry-leading companies: Kweichow Moutai (600519.SH), China Merchants Bank (600036.SH), CATL (300750.SZ), BYD (002594.SZ), LONGi Green Energy (601012.SH), ZTE (000063.SZ) and Mindray (300760.SZ). These stocks span consumer goods, finance, new energy, telecommunications, and healthcare, representing leading companies with strong market influence, stable performance, and innovative potential. For CSI, we conducted a one-year simulation from January 2 to December 31, 2024, using the same initial capital setup, while excluding weekends and official trading holidays. The data set is sourced from Baostock and includes daily records such as the opening price, closing price, highest price, lowest price, trading volume, and multiple technical indicators. This comprehensive data set provides a solid foundation for analyzing market trends and investment strategies within the CSI framework.
The training and testing periods for StockFormer (Gao et al., 2023), TimesNet (Wu et al., 2023), StockMixer (Fan and Shen, 2024), and our system is shown in Table˜7. StockFormer, TimesNet, and StockMixer require longer training periods, and their training data sets span a broader time range compared to the other models. In contrast, our system does not require training and relies solely on historical stock data to make trading decisions.
Strategy NASDAQ CSI
Training Testing Training Testing
Start End Start End Start End Start End
Buy & Hold w/o 25/8/2024 30/8/2024 3/9/2024 29/10/2024 w/o 21/12/2023 29/12/2023 2/1/2024 31/12/2024
SMA w/o 1/8/2024 30/8/2024 3/9/2024 29/10/2024 w/o 1/12/2023 29/12/2023 2/1/2024 31/12/2024
ZMR w/o 10/8/2024 30/8/2024 3/9/2024 29/10/2024 w/o 8/12/2023 29/12/2023 2/1/2024 31/12/2024
MACD w/o 4/8/2024 30/8/2024 3/9/2024 29/10/2024 w/o 4/12/2023 29/12/2023 2/1/2024 31/12/2024
StockFormer w/ 3/7/2023 30/8/2024 3/9/2024 29/10/2024 w/ 2/1/2020 29/12/2023 2/1/2024 31/12/2024
TimesNet w/ 3/7/2023 30/8/2024 3/9/2024 29/10/2024 w/ 2/1/2020 29/12/2023 2/1/2024 31/12/2024
StockMixer w/ 3/7/2023 30/8/2024 3/9/2024 29/10/2024 w/ 2/1/2020 29/12/2023 2/1/2024 31/12/2024
Ours w/o 16/8/2024 30/8/2024 3/9/2024 29/10/2024 w/o 1/12/2023 29/12/2023 2/1/2024 31/12/2024
Table 7: Training and Testing Periods for Various Models on NASDAQ and CSI.
We also select two stocks from the CSI dataset, each representing a different growth pattern, for simulation. As shown in Table˜8, our method consistently outperforms the average trend of the selected stocks.
Trend LLMs TR $\uparrow$ SR $\uparrow$
Average Min Max
-5.86 -13.77 -0.21 GPT4o -0.50 -0.021
Qwen 2.01 0.207
LlaMa3 1.44 0.122
11.37 -4.25 26.05 GPT4o 20.61 0.465
Qwen 16.46 0.420
LlaMa3 4.14 0.167
Table 8: Performance on CSI Stocks with Different Growth Patterns. Our method consistently outperforms the average trend of both selected stocks.
Appendix C Prompt Design Across Modalities
The prompts used for stock analysis with text-only, vision-only, and combined inputs are shown below. Additional prompts for others can be found at: https://github.com/wekjsdvnm/Agent-Trading-Arena.
Text-only Prompt Design Instructions: - You are a player participating in a Simulated Stock Trading Challenge where the prices of stocks are determined mainly by the bids made by the participants in the challenge. - You will be given information related to specific stocks, the whole market, and your existing investments, which is used to analyze the current stock market and investment situation. - Your overall objective is to make as much profit as possible. Stock information: - A: - The closing prices in the past 10 days are: [437.80, 459.44, 465.25, 490.38, 501.034, 511.65, 511.72, 511.79, 511.78, 511.79] - Dividend per share: 22 - Current price change: +0.00%, Current price: 511.79 - Intraday High: 511.79 - Intraday Low: 511.79 - Intraday Mean: 511.79 - B: - The closing prices in the past 10 days are: [460.75, 465.80, 493.27, 502.06, 502.49, 497.32, 486.28, 468.01, 480.61, 480.61] - Dividend per share: 23 - Current price change: +0.00%, Current price: 480.61 - Intraday High: 480.61 - Intraday Low: 480.61 - Intraday Mean: 480.61 - C: - The closing prices in the past 10 days are: [455.90, 440.532, 424.91, 419.75, 420.48, 421.31, 420.12, 421.09, 435.33, 435.33] - Dividend per share: 25 - Current price change: +-0.00%, Current price: 435.33 - Intraday High: 435.33 - Intraday Low: 435.33 - Intraday Mean: 435.33 Market information: Current market index change: 0.58% Gossip from other people: - Exciting developments in the stock market are stirring, with whispers suggesting a potential merger between companies holding Stock A and B. This merger is rumored to lead to significant synergies, possibly resulting in a stock price surge for both A and B. Consequently, many traders anticipate a rise in demand for these stocks, suggesting a timely investment could yield remarkable returns. Meanwhile, analysts rave about Stock C’s steady position and attractive dividends, marking it as a fortress stock that continues to deliver…… Existing Investments: Your total portfolio balance is 134807.00, you are holding the following stocks: - B: - You have held 162 shares of this stock - the total portfolio value is 73839.6 the total capital gain is 5.00% PROFIT - The prices in the past 10 days are: prices: [460.75, 465.80, 493.27, 502.06, 502.49, 497.32, 486.28, 468.01, 480.61, 480.61] - current price change: +0.00%, current price: 480.61, cost price: 455.8 - A: - You have held 130 shares of this stock - the total portfolio value is 60967.4 the total capital gain is 6.00% PROFIT - The prices in the past 10 days are: prices: [437.80, 459.44, 465.25, 490.38, 501.034, 511.65, 511.72, 511.79, 511.78, 511.79] - current price change: +0.00%, current price: 511.79, cost price: 468.98 Investment strategy: try to maximize profit. Task: - Based on (but not limited to) the following information, you need to summarize and conclude the 3 most important and valuable analysis results from the above information. The analysis results must fully fit the focus of the investment strategy, and each result should describe the relationship between its content and investment strategy. In the next stage, the analysis results will be referred for the decision to buy or sell: - Stock information - Market information - Your existing investments - Your investment strategy - Gossip from other people (Gossip may be real or fake news.) Output the response to the prompt above in JSON. Each analysis result should be started with "-", and ended with a line break. Please only provide the response in the following format: {"output": "The analysis results: [analysis results]"}
Visual-only Prompt Design Instructions: - You are a player participating in a Simulated Stock Trading Challenge where the prices of stocks are determined mainly by the bids made by the participants in the challenge. - You will be given information related to specific stocks, the whole market, and your existing investments, which is used to analyze the current stock market and investment situation. - Your overall objective is to make as much profit as possible. Stock information: [Uncaptioned image] (a) Stock A Price [Uncaptioned image] (b) Stock B Price [Uncaptioned image] (c) Stock C Price [Uncaptioned image] (d) Buy & Sell Assert [Uncaptioned image] (e) Trading Record Market information: Current market index change: 0.58% Gossip from other people: - Exciting developments in the stock market are stirring, with whispers suggesting a potential merger between companies holding Stock A and B. This merger is rumored to lead to significant synergies, possibly resulting in a stock price surge for both A and B. Consequently, many traders anticipate a rise in demand for these stocks, suggesting a timely investment could yield remarkable returns. Meanwhile, analysts rave about Stock C’s steady position and attractive dividends, marking it as a fortress stock that continues to deliver consistent profits with minimal risk. As traders weigh their options, strategic portfolio rebalancing tips towards these dividends-rich stocks seem to be the talk of the town. However, with the volatile nature of stock rumors, it remains essential to tread carefully and keep an eye on reliable sources for confirmation. Will this buzz hold substance, or is it merely hot air? Only time will reveal the truth behind the whispers, leaving the market in suspense. Existing Investments: Your total portfolio balance is 134807.00, you are holding the following stocks: - C: - You have held 162 shares of this stock - the total portfolio value is 73839.6 the total capital gain is 5.00% PROFIT - A: - You have held 130 shares of this stock - the total portfolio value is 60967.4 the total capital gain is 6.00% PROFIT Investment strategy: try to maximize profit. Task: - Based on (but not limited to) the following information, you need to summarize and conclude the 3 most important and valuable analysis results from the above information. The analysis results must fully fit the focus of the investment strategy, and each result should describe the relationship between its content and investment strategy. In the next stage, the analysis results will be referred for the decision to buy or sell: - Stock information - Market information - Your existing investments - Your investment strategy - Gossip from other people (Gossip may be real or fake news.) - The provided stock information visualizations. Output the response to the prompt above in JSON. Each analysis result should be started with "-", and ended with a line break. Please only provide the response in the following format: {"output": "The analysis results: [analysis results]"}
Combined-Input Prompt Design Instructions: - You are a player participating in a Simulated Stock Trading Challenge where the prices of stocks are determined mainly by the bids made by the participants in the challenge. - You will be given information related to specific stocks, the whole market, and your existing…… Stock information: - A: - The closing prices in the past 10 days are: [437.80, 459.44, 465.25, 490.38, 501.034, 511.65, 511.72, 511.79, 511.78, 511.79] - Dividend per share: 22 - Current price change: +0.00%, Current price: 511.79 - Intraday High: 511.79 - Intraday Low: 511.79 - Intraday Mean: 511.79 - B: - The closing prices in the past 10 days are: [460.75, 465.80, 493.27, 502.06, 502.49, 497.32, 486.28, 468.01, 480.61, 480.61] - Dividend per share: 23 - Current price change: +0.00%, Current price: 480.61 - Intraday High: 480.61 - Intraday Low: 480.61 - Intraday Mean: 480.61 - C: - The closing prices in the past 10 days are: [455.90, 440.532, 424.91, 419.75, 420.48, 421.31, 420.12, 421.09, 435.33, 435.33] - Dividend per share: 25 - Current price change: +-0.00%, Current price: 435.33 - Intraday High: 435.33 - Intraday Low: 435.33 - Intraday Mean: 435.33 [Uncaptioned image] (a) Stock A Price [Uncaptioned image] (b) Stock B Price [Uncaptioned image] (c) Stock C Price [Uncaptioned image] (d) Buy & Sell Assert [Uncaptioned image] (e) Trading Record Market information: Current market index change: 0.58% Gossip from other people: - Exciting developments in the stock market are stirring, with whispers suggesting a potential merger between companies holding Stock A and B. This merger is rumored to lead to significant synergies…… Existing Investments: Your total portfolio balance is 134807.00, you are holding the following stocks: - B: - You have held 162 shares of this stock - the total portfolio value is 73839.6 the total capital gain is 5.00% PROFIT - The prices in the past 10 days are: prices: [460.75, 465.80, 493.27, 502.06, 502.49, 497.32, 486.28, 468.01, 480.61, 480.61] - current price change: +0.00%, current price: 480.61, cost price: 455.8 - A: - You have held 130 shares of this stock - the total portfolio value is 60967.4 the total capital gain is 6.00% PROFIT - The prices in the past 10 days are: prices: [437.80, 459.44, 465.25, 490.38, 501.034, 511.65, 511.72, 511.79, 511.78, 511.79] - current price change: +0.00%, current price: 511.79, cost price: 468.98 Investment strategy: try to maximize profit. Task: - Based on (but not limited to) the following information, …… describe the relationship between its content and investment strategy. In the next stage, the analysis results will be referred for the decision to buy or sell: - Stock information - Market information - Your existing investments - Your investment strategy - Gossip from other people (Gossip may be real or fake news.) - The provided stock information visualizations. Output the response to the prompt above in JSON. Each analysis result should be started with "-",…… Please only provide the response in the following format: {"output": "The analysis results: [analysis results]"}
Appendix D Simulation Process
The experiments involved several LLMs, including LLaMa-3 (Dubey et al., 2024), GPT-4o (Hurst et al., 2024), GPT-4o-mini (Hurst et al., 2024), DeepSeek (Liu et al., 2024a), Qwen-2.5 (Yang et al., 2024), Qwen-VL-32k (Wang et al., 2024b), Qwen-VL-128kk (Wang et al., 2024b), and Gemini-1.5 (Reid et al., 2024), corresponding to the models Meta-LLaMa-3-70B-Instruct, gpt-4o-2024-08-06, gpt-4o-mini-2024-07-18, DeepSeek-chat, Qwen2.5-72B-Instruct, qwen-vl-max, qwen-vl-max-latest, and Gemini-1.5-pro.
Appendix E Visualization Input
Fig.˜9 illustrates the system’s input prompts and corresponding outputs during the strategy update process. The input prompts consist of both textual and visual components, including daily K-line charts, transaction histories, and agent trading volumes, all of which inform the strategy update.
Appendix F Simulation Process
In the Agent Trading Arena, the simulation process unfolds as follows: First, rumors are generated in the chat pool based on the previous day’s stock market analysis. Next, historical stock data is analyzed, followed by decision-making and execution. Short-term memory is formed through interactions with the environment. Finally, the system evaluates this memory, updates the strategy, and consolidates it into long-term memory. This entire process is illustrated in Fig.˜10.
@@ Appendix F Simulation Process
Impact of Modality on LLM Competitiveness.
We employed a relative evaluation method for this experiment. The first and second agents used various LLMs in textual and visual settings, respectively, while the remaining agents were based on LLaMa-3 (Dubey et al., 2024) as the baseline. This setup aimed to explore the impact of different agents and modalities on LLM performance. The results are shown in Table˜9. The findings indicate that DeepSeek (Liu et al., 2024a) exhibited stronger competitive performance across different LLM environments, suggesting unique strengths or optimizations that enable it to adapt more effectively to the task’s complexities.
LLMs T V TR $\uparrow$ SR $\uparrow$
LLaMa-3 100.00 100.00
DeepSeek + 35.03 + 45.63
Gemini-1.5 + 10.99 + 14.93
+ 11.49 + 29.47
GPT-4o + 10.39 + 28.29
+ 17.18 + 40.77